OpenStudy (anonymous):

You deposit $100 at the end of each month into an annuity that pays 6.5% compounded monthly. What is the value of the annuity after 5 years? Round to the nearest cent

OpenStudy (anonymous):

anybody?

OpenStudy (anonymous):

\[Final = (Deposit) \frac{ (1+i)^n - 1 }{ i }\] where, i = interest and n = number of times compounded. \[i = 0.065\]\[n = 12 \times 5 = 60\] \[Final = 100\frac{ (1+ 0.065)^{60}-1 }{ 0.065 } \approx 65,768.99\]

OpenStudy (anonymous):

Uhhhhh... where did that come from? It's 7067.40